# biblio.bib Metaheuristics and Large Language Models Join Forces: Towards an Integrated Optimization Approach

### v1

@@ 5.2.2 Dimension 2 of the Evaluation Framework: Alternative Techniques for Guiding MHs
We believe that, although BRKGA+LLM’s results are not favorable compared to those obtained with the help of irace in the context of two out of four problem instances, our approach benefits from a significantly reduced computational effort. This still holds when the computation time required to extract five metric values to build the prompts is taken into account. Additionally, with model improvements or prompt adjustments—for example, the LLM currently, by its own internal decision, produces values divisible by 0.05, leading to a loss of precision compared to irace—greater precision could enable the LLM to improve in the future.
5.2.3 Dimension 3 of the Evaluation Framework: Prompt Quality
In the final dimension, we transition from a numerical analysis—as conducted in the previous two dimensions—to a more interpretative approach, where we discuss the complex process of designing a well-working prompt tailored to the considered optimization problem. To accomplish this, we first examine the five selected metrics already introduced in Section 4.1. This is crucial because every additional metric enlarges the prompt. And the larger the prompt, the smaller the maximum network size that—due to the limited size of the context window—can be passed to the LLM. Also, larger prompts are associated with increased financial costs. Hence, it is essential to determine whether similar or even superior results can be achieved using less information. To investigate this, we perform two experiments: the first assesses the correlation between each pair of metrics, and the second removes information from the prompt to analyze the impact on the LLM results.
Correlation between metrics. For this analysis, we utilize a matrix of plots provided in Figure 5 concerning the soc-hamsterster instance for which BRKGA+LLM significantly outperformed BRKGA; see Table 5. Hereby, the plots in the upper triangle of the matrix are scatter plots that display the values of all pairs of metrics. For example, the second plot in the first matrix row shows the scatter plot concerning the values of metrics out-degree (x-axis) and in-degree (y-axis). In contrast, the plots in the lower triangle are kernel density estimation (KDE) plots for each pair of metrics. KDE plots provide a smoothed representation of the underlying data distribution, aiding in the identification of patterns such as clusters, outliers, and non-linear relationships. Lastly, the diagonal showcases univariate KDE plots for each variable, analogous to a histogram, depicting the distribution of each variable independently. Note that a corresponding graphic concerning the other problem instances is provided in the GitHub repository, whose link can be found in Section 5.1.3. The following observations can be made based on Figure 5:
@@ 5.2.3 Dimension 3 of the Evaluation Framework: Prompt Quality
Upon examining each metric, we conclude that all are potentially relevant and contribute to the outcome. However, it is still possible that adjusting the set of metrics by removing certain metrics or adding alternative ones could lead to even better results.171717Financial limitations prevent from conducting extensive experiments and from studying every aspect of the prompt, especially for large-size, challenging scenarios; hence, careful consideration is essential.
@@
Differences in node selection. Finally, we wanted to study how the use of LLM output leads to the selection of different nodes for solutions produced by BRKGA+LLM in comparison to BRKGA. For this purpose, Figure 6 shows the node probabilities computed based on the alpha and beta values (black line) in relation to the (normalized) values of the five metrics, exemplary in the context of the synthetic graph 0.2-0.0-0.3-0.5. The x-axis ranges over all 500 graph nodes, ordered by a non-increasing LLM-probability. Moreover, by means of horizontal lines the graphic marks the nodes chosen for the best BRKGA solution (dotted), the best BRKGA+LLM solution (solid), and their intersection (dashed). In the following, we point out three specific cases highlighted as (a), (b), and (c) in Figure 6:
@@
5.3 Visual Comparative Analysis
Numerical analysis can fall short of capturing the full complexity of a metaheuristics’ search process due to its stochastic nature. Visual tools have emerged in recent years to address this limitation. They were developed to provide a more comprehensive understanding and additional insight. One such tool is STNWeb [CHACONSARTORI2023100558], which generates directed graphs from algorithm trajectories to visualize how these algorithms navigate the search space. This allows to compare and justify the performance of different algorithms. A visual analysis is presented in this section to understand better the advantages of our BRKGA+LLM approach over BRKGA and BRKGA+FC.
Figure 7 shows a STNWeb plot displaying 10 runs of each BRKGA+LLM, BRKGA, and BRKGA+FC when applied to the soc-hamsterster instance. The following analysis highlights the insights that can be obtained from this visualization. However, first of all, let us explain the technical elements of the plot:
@@ 5.3 Visual Comparative Analysis
Several interesting observations can be made based on Figure 7. First of all, only BRKGA+LLM can find best solutions (see the two red dots). Moreover, the two best solutions found by BRKGA+LLM are rather different from each other, as they are found in different areas of the search space. The three algorithms seem to be attracted by different areas of the search space. Moreover, while BRKGA and BRKGA+FC clearly converge to solutions that are close to each other, this is not so much the case for BRKGA+LLM, which does not show a clear convergence behavior towards a single area of the search space. Finally, note that the search trajectories of BRKGA are much shorter than those of the two hybrid approaches. Additional STN plots concerning other problem instances can be found in the repository whose link was provided in Section 5.1.3.
Following the empirical study of our prompt’s design and quality (Section 5.2) and following a visual analysis, we can conclude that our proposed hybridization achieves its intended objective: showing that LLMs can generate heuristic information that can be used to improve the search process of a metaheuristic. Interestingly, our approach has also outperformed an alternative hybridization scheme that used heuristic information produced by a hand-crafted, specifically trained, graph neural network from [9909110]. However, successfully integrating LLMs into MHs involves addressing several critical issues and open questions, detailed in the next section.
6 Discussion and open questions
@@
Are LLMs merely stochastic parrots or black boxes capable of reasoning? In the influential paper by Bender et al. [10.1145/3442188.3445922], the authors raise concerns about the risks of using LLMs and question their necessity. They argue that LLMs are trained on vast amounts of data based on probability distributions but lack any reference to meaning, earning them the label of stochastic parrots. However, significant progress has been made since the paper’s publication in 2021. LLMs have improved their capabilities, and it is questionable if they can still be called “stochastic parrots.” Instead, they might be seen as black boxes that can reason within a certain context. Our research demonstrates the utility of LLMs in one such context of reasoning. Moreover, several other studies have confirmed the abil of LLMs to reason in other specific contexts [ahn2024large, mirchandani2023large]. On the other hand, other —more sensitive— contexts such as legal or moral reasoning [ALMEIDA2024104145] need more thoughtful human oversight, since a careless integration of LLMs could potentially have an unpredictably disruptive effect. Thus, we suggest approaching LLMs with caution and letting the experiments speak for themselves. After all, technology can improve rapidly, and it is essential to avoid hasty dismissal or overhyping LLMs as a silver bullet.
@@
Are private LLMs the sole providers of superior outcomes? In the past, it would have been accurate to affirm this, as models such as GPT-4 and Claude-3-Opus were recognized for their top-tier response quality. In fact, our research demonstrates that Claude-3-Opus outperforms its competitors. However, the scene is evolving rapidly. Recent models, such as Cohere’s Command-R+, Mistral’s Mistral models, and Meta’s Llama 3 (which we did not incorporate in our study, because its context limit is very low: ), are delivering results comparable to some of the before-mentioned models (e.g., [davis2024prompting, liu2023llm360, chen2024chatgpts]). These models also come with open licenses, although with varying levels of permissiveness. Nonetheless, it is important to note that these models are all products of private entities with substantial financial resources. The prospect of a public entity or small business independently creating an LLM from scratch appears remote, largely due to the computational demands and associated costs. This could potentially pose a risk to the transparency of their design processes [harandizadeh2024risk].
@@
What are the primary obstacles to adopting the strategy proposed in our research? We identify two significant hurdles: cost constraints and technical limitations. While leveraging LLMs as software-as-service can alleviate the need for in-house cluster training, a substantial financial burden is incurred by processing large volumes of tokens. Our investigation revealed that even with moderate-sized graph instances of around nodes, we quickly reach the token limit of what the most permissive LLM can handle within its context window (see subsection 5.1.2). Having said that, recent studies (see, for example, [munkhdalai2024leave]) have proposed the possibility of “infinite” context windows. Additionally, Google’s Gemini Pro 1.5 model boasts an impressive -token context window limit [geminiteam2024gemini]. However, to successfully apply a strategy similar to ours in the context of massive graphs—or, more generally, in the context of large-scale problem instances—these two limitations must be significantly mitigated. Alternatively, a novel prompt strategy could be developed to detect patterns in metrics while reducing token counts, potentially through prompt compression [jiang2023llmlingua] or relevant node metric filtering. Neither of these alternatives has been investigated in our research.
@@
As researchers, what other aspects of LLMs should we be cautious about? Earlier, we touched on the issue of LLMs’ capability for reasoning in certain contexts. However, many argue that they are incapable of reasoning altogether. This discrepancy stems from the ambiguity surrounding the term “reasoning”. In our context, reasoning refers to the ability to infer and identify useful patterns in metrics associated with each node of a graph. In particular, we demonstrated that LLMs do not produce random or uninformative values but rather follow the given instructions. However, if we had left the concept of “reasoning” or “pattern discovery” too vague, our research would be susceptible to misinterpretation. Therefore, we recommend that researchers be aware of the underlying philosophical discussions surrounding this technology when working with LLMs. After all, technological advancements can shape the way we express ourselves. A good starting point may be to engage with the works of Floridi (e.g., [Floridi2023-FLOAAA-4, Floridi2024-jv]). Particularly since the boundaries of LLMs’ capabilities remain undefined and are a subject of ongoing debate.
@@
Are there ways to improve the integration between MHs and LLMs? As discussed in Section 2.2, there are already a few hybridization approaches. These include creating new MHs by leveraging LLMs to generate code and employing LLMs as solvers for optimization problems described in natural language. In contrast, our approach utilizes LLMs as pattern detectors for complex instances. These patterns are then used to bias the search process of the metaheuristic. However, an intriguing integration could involve combining all these hybridization techniques within a single software framework. We believe that these approaches are not mutually exclusive but rather complementary. By unifying them, we may unlock even greater contributions to the field of MHs. One potential approach to achieving this integration could be through the use of agents [xi2023rise], which would be responsible for orchestrating the three hybridization methods. Note that agents are currently a topic of significant interest in the LLM community [guo2024large].
7 Conclusion
@@
This paper showcased the potential of leveraging Large Language Models (LLMs) as pattern search engines to enhance metaheuristics (MH) by integrating the information they provide. We demonstrate the effectiveness of this approach on a combinatorial optimization problem in the realm of social networks. An important aspect of our work is prompt engineering. In fact, useful LLM answers are only obtained with well-designed prompts. In the context of the considered social networks problem, the LLM output is used to compute a probability for each node of the input graph to belong to an optimal solution. These node probabilities are then used to bias the search process of a biased random key genetic algorithm (BRKGA). We could show that our hybrid approach outperforms both the pure BRKGA and the state-of-the-art BRKGA variant, whose search process is biased by the output of a hand-designed (and trained) graph neural network model.
@@
Acknowledgements
C. Chacón, C. Blum and F. Bistaffa were supported by grants TED2021-129319B-I00 and PID2022-136787NB-I00 funded by MCIN/AEI/10.13039/501100011033.

### v2

@@ V-B2 Dimension 2 of the Evaluation Framework: Alternative Techniques for Guiding MHs
We believe that, although BRKGA+LLM’s results are not favorable compared to those obtained with the help of irace in the context of two out of four problem instances, our approach benefits from a significantly reduced computational effort (see TableVII). This still holds when the computation time required to extract five metric values to build the prompts is taken into account. Additionally, with model improvements or prompt adjustments—for example, the LLM currently, by its own internal decision, produces values divisible by 0.05, leading to a loss of precision compared to irace—greater precision could enable the LLM to improve in the future.
V-B3 Dimension 3 of the Evaluation Framework: Prompt Quality
In the final dimension, we transition from a numerical analysis—as conducted in the previous two dimensions—to a more interpretative approach, where we discuss the complex process of designing a well-working prompt tailored to the considered optimization problem. To accomplish this, we first examine the five selected metrics already introduced in Section IV-A. This is crucial because every additional metric enlarges the prompt. And the larger the prompt, the smaller the maximum network size that—due to the limited size of the context window—can be passed to the LLM. Also, larger prompts are associated with increased financial costs. Hence, it is essential to determine whether similar or even superior results can be achieved using less information. To investigate this, we perform two experiments: the first assesses the correlation between each pair of metrics, and the second removes information from the prompt to analyze the impact on the LLM results.
Correlation between metrics. For this analysis, we utilize a matrix of plots provided in Figure V-B3 concerning the soc-hamsterster instance for which BRKGA+LLM significantly outperformed BRKGA; see Table V. Hereby, the plots in the upper triangle of the matrix are scatter plots that display the values of all pairs of metrics. For example, the second plot in the first matrix row shows the scatter plot concerning the values of metrics out-degree (x-axis) and in-degree (y-axis). In contrast, the plots in the lower triangle are kernel density estimation (KDE) plots for each pair of metrics. KDE plots provide a smoothed representation of the underlying data distribution, aiding in the identification of patterns such as clusters, outliers, and non-linear relationships. Lastly, the diagonal showcases univariate KDE plots for each variable, analogous to a histogram, depicting the distribution of each variable independently. Note that a corresponding graphic concerning the other problem instances is provided in the GitHub repository, whose link can be found in Section I-B. The following observations can be made based on Figure V-B3:
@@ V-B3 Dimension 3 of the Evaluation Framework: Prompt Quality
Upon examining each metric, we conclude that all are potentially relevant and contribute to the outcome. However, it is still possible that adjusting the set of metrics by removing certain metrics or adding alternative ones could lead to even better results.202020Financial limitations prevent from conducting extensive experiments and from studying every aspect of the prompt, especially for large-size, challenging scenarios; hence, careful consideration is essential.
[h!](topskip=0pt, botskip=0pt, midskip=0pt)[width=1]figures/metrics.pdf Correlations between all pairs of the five considered metrics concerning the soc-hamsterster network.
@@
Differences in node selection. Finally, we wanted to study how the use of LLM output leads to the selection of different nodes for solutions produced by BRKGA+LLM in comparison to BRKGA. For this purpose, Figure V-B3 shows the node probabilities computed based on the alpha and beta values (black line) in relation to the (normalized) values of the five metrics, exemplary in the context of the synthetic graph 0.2-0.0-0.3-0.5. The x-axis ranges over all 500 graph nodes, ordered by a non-increasing LLM-probability. Moreover, by means of horizontal lines the graphic marks the nodes chosen for the best BRKGA solution (dotted), the best BRKGA+LLM solution (solid), and their intersection (dashed). In the following, we point out three specific cases highlighted as (a), (b), and (c) in Figure V-B3:
1.
@@
[h!](topskip=0pt, botskip=0pt, midskip=0pt)[width=1]figures/analysis-metrics.pdf Analysis of the probabilities computed based on the alpha and beta values (black line) in relation to the (normalized) values of the five metrics. The x-axis ranges of all 500 nodes of the synthetic graph 0.2-0.0-0.3-0.5 ordered by a non-increasing LLM-probability. Moreover, the graphic marks the nodes chosen for the best BRKGA solution, the best BRKGA+LLM solution, and their intersection.
V-C Visual Comparative Analysis
Numerical analysis can fall short of capturing the full complexity of a metaheuristics’ search process due to its stochastic nature. Visual tools have emerged in recent years to address this limitation. They were developed to provide a more comprehensive understanding and additional insight. One such tool is STNWeb [47], which generates directed graphs from algorithm trajectories to visualize how these algorithms navigate the search space. This allows to compare and justify the performance of different algorithms. A visual analysis is presented in this section to understand better the advantages of our BRKGA+LLM approach over BRKGA and BRKGA+FC.
[h](topskip=0pt, botskip=0pt, midskip=0pt)[width=1]figures/stn.pdf STNWeb-generated plot comparing the trajectories of BRKGA (cyan), BRKGA+FC (magenta) and BRKGA+LLM (green) over 10 runs on the soc-hamsterster instance (with and ). This plot was generated using the so-called agglomerative clustering partitioning method available in STNWeb, with the number of clusters set to approximately 20% of the total, allowing for a visualization focusing on the essential characteristics.
Figure V-C shows a STNWeb plot displaying 10 runs of each BRKGA+LLM, BRKGA, and BRKGA+FC when applied to the soc-hamsterster instance. The following analysis highlights the insights that can be obtained from this visualization. However, first of all, let us explain the technical elements of the plot:
@@ V-C Visual Comparative Analysis
Several interesting observations can be made based on Figure V-C. First of all, only BRKGA+LLM can find best solutions (see the two red dots). Moreover, the two best solutions found by BRKGA+LLM are rather different from each other, as they are found in different areas of the search space. The three algorithms seem to be attracted by different areas of the search space. Moreover, while BRKGA and BRKGA+FC clearly converge to solutions that are close to each other, this is not so much the case for BRKGA+LLM, which does not show a clear convergence behavior towards a single area of the search space. Finally, note that the search trajectories of BRKGA are much shorter than those of the two hybrid approaches. Additional STN plots concerning other problem instances can be found in the repository whose link was provided in Section I-B.
Following the empirical study of our prompt’s design and quality (Section V-B) and following a visual analysis, we can conclude that our proposed hybridization achieves its intended objective: showing that LLMs can generate heuristic information that can be used to improve the search process of a metaheuristic. Interestingly, our approach has also outperformed an alternative hybridization scheme that used heuristic information produced by a hand-crafted, specifically trained, graph neural network from [19]. However, successfully integrating LLMs into MHs involves addressing several critical issues and open questions, detailed in the next section.
VI Discussion and open questions
@@
Are LLMs merely stochastic parrots or black boxes capable of reasoning? In the influential paper by Bender et al. [48], the authors raise concerns about the risks of using LLMs and question their necessity. They argue that LLMs are trained on vast amounts of data based on probability distributions but lack any reference to meaning, earning them the label of stochastic parrots. However, significant progress has been made since the paper’s publication in 2021. LLMs have improved their capabilities, and it is questionable if they can still be called “stochastic parrots.” Instead, they might be seen as black boxes that can reason within a certain context. Our research demonstrates the utility of LLMs in one such context of reasoning. Moreover, several other studies have confirmed the abil of LLMs to reason in other specific contexts [5, 32]. On the other hand, other —more sensitive— contexts such as legal or moral reasoning [4] need more thoughtful human oversight, since a careless integration of LLMs could potentially have an unpredictably disruptive effect. Thus, we suggest approaching LLMs with caution and letting the experiments speak for themselves. After all, technology can improve rapidly, and it is essential to avoid hasty dismissal or overhyping LLMs as a silver bullet.
@@
Are private LLMs the sole providers of superior outcomes? In the past, it would have been accurate to affirm this, as models such as GPT-4 and Claude-3-Opus were recognized for their top-tier response quality. In fact, our research demonstrates that Claude-3-Opus outperforms its competitors. However, the scene is evolving rapidly. Recent models, such as Cohere’s Command-R+, Mistral’s Mistral models, and Meta’s Llama 3 (which we did not incorporate in our study, because its context limit is very low: ), are delivering results comparable to some of the before-mentioned models (e.g., [49, 50, 51]). These models also come with open licenses, although with varying levels of permissiveness. Nonetheless, it is important to note that these models are all products of private entities with substantial financial resources. The prospect of a public entity or small business independently creating an LLM from scratch appears remote, largely due to the computational demands and associated costs. This could potentially pose a risk to the transparency of their design processes [52].
@@
What are the primary obstacles to adopting the strategy proposed in our research? We identify two significant hurdles: cost constraints and technical limitations. While leveraging LLMs as software-as-service can alleviate the need for in-house cluster training, a substantial financial burden is incurred by processing large volumes of tokens. Our investigation revealed that even with moderate-sized graph instances of around nodes, we quickly reach the token limit of what the most permissive LLM can handle within its context window (see subsection V-A2). Having said that, recent studies (see, for example, [53]) have proposed the possibility of “infinite” context windows. Additionally, Google’s Gemini Pro 1.5 model boasts an impressive -token context window limit [54]. However, to successfully apply a strategy similar to ours in the context of massive graphs—or, more generally, in the context of large-scale problem instances—these two limitations must be significantly mitigated. Alternatively, a novel prompt strategy could be developed to detect patterns in metrics while reducing token counts, potentially through prompt compression [55] or relevant node metric filtering. Neither of these alternatives has been investigated in our research.
@@
As researchers, what other aspects of LLMs should we be cautious about? Earlier, we touched on the issue of LLMs’ capability for reasoning in certain contexts. However, many argue that they are incapable of reasoning altogether. This discrepancy stems from the ambiguity surrounding the term “reasoning”. In our context, reasoning refers to the ability to infer and identify useful patterns in metrics associated with each node of a graph. In particular, we demonstrated that LLMs do not produce random or uninformative values but rather follow the given instructions. However, if we had left the concept of “reasoning” or “pattern discovery” too vague, our research would be susceptible to misinterpretation. Therefore, we recommend that researchers be aware of the underlying philosophical discussions surrounding this technology when working with LLMs. After all, technological advancements can shape the way we express ourselves. A good starting point may be to engage with the works of Floridi (e.g., [56, 33]). Particularly since the boundaries of LLMs’ capabilities remain undefined and are a subject of ongoing debate.
@@
Are there ways to improve the integration between MHs and LLMs? As discussed in Section II-B, there are already a few hybridization approaches. These include creating new MHs by leveraging LLMs to generate code and employing LLMs as solvers for optimization problems described in natural language. In contrast, our approach utilizes LLMs as pattern detectors for complex instances. These patterns are then used to bias the search process of the metaheuristic. However, an intriguing integration could involve combining all these hybridization techniques within a single software framework. We believe that these approaches are not mutually exclusive but rather complementary. By unifying them, we may unlock even greater contributions to the field of MHs. One potential approach to achieving this integration could be through the use of agents [57], which would be responsible for orchestrating the three hybridization methods. Note that agents are currently a topic of significant interest in the LLM community [31].
VII Conclusion
@@
This paper showcased the potential of leveraging Large Language Models (LLMs) as pattern search engines to enhance metaheuristics (MH) by integrating the information they provide. We demonstrate the effectiveness of this approach on a combinatorial optimization problem in the realm of social networks. An important aspect of our work is prompt engineering. In fact, useful LLM answers are only obtained with well-designed prompts. In the context of the considered social networks problem, the LLM output is used to compute a probability for each node of the input graph to belong to an optimal solution. These node probabilities are then used to bias the search process of a biased random key genetic algorithm (BRKGA). We could show that our hybrid approach outperforms both the pure BRKGA and the state-of-the-art BRKGA variant, whose search process is biased by the output of a hand-designed (and trained) graph neural network model. To address this, we created a tool named OptiPattern (LLM-Powered Pattern Recognition for Combinatorial Optimization), which implements this hybrid method and is available at: https://github.com/camilochs/optipattern.
@@
Acknowledgment
The research presented in this paper was supported by grants TED2021-129319B-I00 and PID2022-136787NB-I00 funded by MCIN/AEI/10.13039/ 501100011033.